Grade 12th passMechanics

TWO CARS OF LENGTH 100M ARE MOVING IN THE SAME DIRECTION WITH SPEEDS 15M/S AND 10M/S ON A PARALLEL TRACK . THEN FIND THE TIME OF OVERTAKING OR CROSSING.

swaati
6 Years agoGrade 12th pass

1 Answer

Vikas TU
6 Years ago
Dear student 
I think you should use  the conept o Relative speed here.
Relative speed between Two cars = 15-10 = 5m/sec
So, The second one will overtake tthe first one in 100/5 = 20 sec.
Hope this helps